Continuous increase in frequency

Hi, I am using the openQCM Q-1 to measure protein adsorption in liquid.

My procedure is as follows:
1. Inject buffer at 100 uL/min for 10 min using a syringe pump to fill the fluidic chamber
2. Stop the pump and allow the frequency to reach a stable value under static conditions (20 min)
3. After 20 min, inject protein sample (in the same buffer) as in Step 1
4. Stop the pump and allow the frequency to reach a new stable value
5. Calculate frequency shift by subtracting the baseline frequency in Step 1 from the frequency in Step 4

Initial experiments were alright, and I was able to obtain negative frequency shifts which indicate protein binding to the surface. However, sometimes the frequency begins to rise continuously halfway through the experiment, rising by up to 4 Hz every 30 min (even though the frequency had been stable at first), and this masks any negative shifts that may be due to protein binding. Temperature fluctuations are very minor, and it does not seem to be caused by bubbles formed.

I had suspected that it was due to mechanical stresses from the O-ring over time, so I tried knocking on the device to release stresses (as mentioned in the manual). This action sometimes helps to restore the frequency value, however, the value is still up to 5 Hz higher (even when tested only with buffer) and releasing the stress during an experiment does not seem to be a good idea.

Have you encountered this problem when performing measurements in liquid involving sample changes? Would appreciate some help in how to solve this problem. Thank you!


  • Hi huijeanlim,
    thank you for your interesting question.
    regarding your procedure, your impression is right. Sometimes QCM instruments have drifts due to o-ring swelling or other mechanical stress on the sensor.
    By knocking on the device fluidic cell, you can release stress due to o-ring swelling. On the other hand, pogo-pin contacts can induce further stress on the sensor.
    In order to solve this point, we design a true hole on the basis of the sensor module (bottom side). By using a hex key of 1.3 mm you can fine-tune the pogo-pin height. In this way, you should solve the drift issue. Anyway, if you will still observe drifts, contact me again, in order to help you.
  • Hi Raffaele, thanks a lot for the reply. Will give these methods a try.
  • You are welcome and thank you too. Please let me know if you will solve!
Sign In or Register to comment.